OpenStudy (anonymous):

What is the solution for these three equations? 1/2 x - 2y = 4 and 4y - x = -8 3x + 2y = 6 and y = -3/2 x - 1 3x - 2y = 6 and 5x + 3y = -9

OpenStudy (anonymous):

1/2 x - 2y = 4 and 4y - x = -8 x=8,y=0 3x + 2y = 6 and y = -3/2 x - 1 I couldn't solve this. 3x - 2y = 6 and 5x + 3y = -9 x=0,y=-3

OpenStudy (anonymous):

thank you! and I already solve it no problem :)